If the sides of a square are lengthened by 8cm, the area becomes 196cm^2. Find the length of a side of the original square.
Answer by richwmiller(17219)

x=side of square
x+8=new side
(x+8)^2=196
|x+8|=14
x+8=-14
x=-22
x+8=14
x=6 cm
we can only use a positive number for the side of a square. so we have 6 cm as the original side of the square